Use the clues to find the value of E.
A x B = 80
A x C = 200
B x D = 36
D x E = 18
C = 100

Found 2 solutions by Edwin McCravy, rothauserc:
Answer by Edwin McCravy(20081)

AB = 80
AC = 200
BD = 36
DE = 18
C = 100 

AC = 200
and since C=100
A(100) = 200
A=200/100=2

AB = 80
and since A=2
2B = 80
 B = 80/2 = 40

BD = 36
and since B=40
40D = 36
D = 36/40 = 9/10

DE = 18
and since D=9/10
(9/10)E = 18
E = 18÷(9/10) = 18×(10/9) = 20

Answer: E=20

Edwin

Answer by rothauserc(4718)

From the given clues,
:
80/B = 200/C = 200/100 = 2/1
:
2B = 80
:
B = 40
:
36/B = 18/E
:
36E = 18B
:
E = B/2 = 40/2 = 20
:
************
E = 20
************
: